(1)Write the slope- intercept form of the equation of the line through (-1,2)that is perpendicular to the line 3x+y=4.

(2) Find the equation of a line through (-2,0) which is perpendicular to the line -5x-15y=11.


(3) find the slope of the line which is parallel to the line x=4/5y +7/4

Answer by Nate(3500)

(1)
m=1/3 *perpendicular is the opposite of the recripricol* point = (-1,2)
y - (ysub1)=m(x - (xsub1))
y - 2 = (1/3)(x + 1)
y-2=(1/3)x+(1/3)
y=(1/3)x+(7/3)
(2)
m=3 point = (-2,0)
y - (ysub1)=m(x - (xsub1))
y=3(x+2)
y=3x+6
(3)
y=(5/4)x+(35/16)
m=(5/4)